Find the extrema of

3x^2-y^2+x^3


1
Expert's answer
2021-11-15T16:25:09-0500
"f(x, y)=3x^2-y^2+x^3"

"f_x=6x+3x^2"

"f_y=-2y"

Find the critical point(s)


"\begin{matrix}
 f_x=0 \\
 f_y=0
\end{matrix}=>\begin{matrix}
 6x+3x^2=0 \\
 -2y=0
\end{matrix}"

"Point(-2,0), Point(0,0)"


"f_{xx}=6+6x"

"f_{xy}=0"

"f_{yy}=-2"

"D=\begin{vmatrix}
 6+6x & 0 \\
 0 & -2
\end{vmatrix}=-12-12x"

"Point(-2,0)"


"f_{xx}=6+6(-2)=-6<0"

"D=-12-12(-2)=12>0"

Then "f(-2, 0)" is a local maximum.


"Point(0,0)"


"f_{xx}=6+6(0)=6>0"

"D=-12-12(0)=-12<0"

Then "Point(0,0)" is a saddle point.



"f(-2, 0)=3(-2)^2-(0)^2+(-2)^3=4"